OpenStudy (anonymous):

Find the cubic function of the form y=ax^3 + bx^2 + cx + d which has a relative maximum point at (0,2) and a point of inflection at (-1,-2). How do I even start this one?

OpenStudy (anonymous):

take the derivative, set it equal zero, and solve

OpenStudy (anonymous):

then \[y''=6ax+2b\] and you want this to be zero when x = -1, so you get \[6a(-1)+2b=0\] \[-6a+2b=0\]

OpenStudy (anonymous):

ok i don't know why i deleted my reply, because it was right so i write it again. you know \[y'=3ax^2+2bx+c=0\] if x = 0 that tells you \[c=0\]

OpenStudy (anonymous):

and we also know that if x = 0, then y = 2 which tells you \[d=2\]

OpenStudy (anonymous):

so we have \[y=ax^3+bx^2+2\]

OpenStudy (anonymous):

and now use that if x = -1, y = -2 to write \[-a+b+2=-2\] \[-a+b=-4\] and solve the system \[-a+b=-4\] \[-6a+2b=0\] for "a" and "b"